OpenStudy (anonymous):

find the solution of the inequality |x - 5| ≥ 3

OpenStudy (p0sitr0n):

take x-5 and 5-x >= 3

OpenStudy (cwrw238):

this is equivalent to x - 5 >= 3 and x - 5 <= -3 solve these 2 inequalities and you have it

OpenStudy (cwrw238):

can you solve x - 5 >= 3 ?

OpenStudy (cwrw238):

do you know what the | | (= ' mod' ) around an expression means ?

OpenStudy (anonymous):

absolute value

OpenStudy (p0sitr0n):

it means x-5 can be either positive or negative. so take (x-5) and -(x-5)

OpenStudy (anonymous):

okay

OpenStudy (cwrw238):

thats why you have 2 possibilities

OpenStudy (cwrw238):

leading to 2 equations

OpenStudy (anonymous):

2 different answers?

OpenStudy (cwrw238):

the answers will be a range of values of x - i should have said 2 inequalities x - 5 >= 3 x >= 8 -now solve the other inequality

OpenStudy (p0sitr0n):

yes, because it depends on wheter the original equation was positive or negtive. its like multiple roots of a quadratic, there can be more than 1

OpenStudy (anonymous):

x <= -8

OpenStudy (anonymous):

@cwrw238

OpenStudy (cwrw238):

no x <= -3 + 5

OpenStudy (cwrw238):

add 5 to both sides of the inequality

OpenStudy (anonymous):

x <= 2?

OpenStudy (cwrw238):

right

OpenStudy (cwrw238):

so combining these 2 results we have x <=2 or x >= 8

OpenStudy (anonymous):

thanks

OpenStudy (cwrw238):

and we can graph it on a number line as follows